T.R | Title | User | Personal Name | Date | Lines |
---|
23.1 | first and last reply for 1996 | FABSIX::M_RIDLON | Eliminate the oblivious | Tue Dec 31 1996 07:21 | 6 |
| Reconfigured the vacuum solenoid valve, gauge, and filter/trap. The
setup matches that on the other speedfams which gets rid of 4 elbows
and over three feet of vacuum hose.
B-shift
|
23.2 | source report | STRATA::TDYER | | Fri Jan 03 1997 09:13 | 161 |
|
+---------------------------+ TM
| d | i | g | i | t | a | l | I N T E R O F F I C E M E M O R A N D U M
+---------------------------+
| S e m i c o n d u c t o r |
+---------------------------+
To: Ellen Mager Date: December 17th, 1996
From: Frank Krupa/Timothy Dyer
Dept: CMP
ENET: SUBPAC::KRUPA,
SCOMAN::TDYER
cc: Stephanie Salem DTN: 225-5737,5072
Dennis Goodwin
Kyle Wooldridge
Kim Conway
MaryAnn Piasecki
Linda Bioseneualt
Ray Tadry
Subject: CMP-V SERIAL #47 SOURCE INSPECTION REPORT
Summary:
Digital arrived on the morning of 12/9/96 to perform a customer source
inspection for serial number 47. It was our immediate impression that the
tool was not ready for the source inspection as was the case with serial
number 24's source inspection. The tool was not fully configured or tested,
see configuration issues listed below. It is the customer's expectation that
prior to customer arrival the polisher has been: 1) FULLY tested by SpeedFam
prior to our arrival; 2) Operating with the latest released version of software
3) The customer's recipe has been robust tested using all aspects of the
software including activating the alarms. In addition our local FSE arrived
several days ahead of us with intention of performing a pre-source inspection
to ensure the system's readiness prior to customer arrival. The FSE was not
able to perform the required testing. As a result of the above not being
completed, the polisher errored for "low down force" on second table during
processing of our first wafers. The actual problem was due to head-1 not
lowering in sufficient time. If the polisher had been tested by SpeedFam and
again by the FSE, this would not have occurred on the customer's time. One
day was lost due to this which should have been avoided. Speedfam needs
to improve their pre-source inspection testing to come up to industry standards.
Once the software was updated, the recipes corrected, and the head-1
problem was corrected, the source acceptance ran without ANY issues. The
500 wafer mechanical marathon ran with out any alarms/errors. Following the
mechanical portion the process acceptance testing began. Speedfam configured
the system for process testing with old generic carriers/plates. As a result
10 wafers and ~2 hours were lost trouble-shooting the problem. After locating
acceptable carriers/plates, testing resumed. A few small offset and ID/OD
adjustments were made and the 50 test wafers were committed. The process
results were acceptable.
On a positive note the system ran without ANY issues once the head-1
problem was corrected. It was also refreshing to have an oxide measurement
system in the same room as the testing was performed and to be the first
customer to use the new spin-rinse dryer in the room(even though we had to
program it).
OPEN ISSUES:
Configuration Issues:
1. A/B boxes not installed
2. Remote terminal and key board not located
3. SECS-II not installed. Installation will require automation frame
separation, replacement of bulkhead panel and installation of
hardware. Due to any possible disturbances, Digital is requesting
an additional 100 wafers to be cycled through the tool after the
SECS-II work is completed..
4. Robot pendent not located
5. Chiller serial number ___________
6. New style door not installed.
Mechanical Issues:
1. Leak at top of fluids manifold
2. Leak near vacuum tank
3. Leak detect not enabled
4. Down force calibrated with load cell and aluminum wafer directly
contacting the carrier film. This procedure causes a 15-20 micron
dip In the carrier films. Procedure needs improvement.
Software Issues:
1. Most current released version not installed, changed from 494a to
494e.
PRODUCTION MANGER WILL SET-UP PROCEDURE WITH SOFTWARE GROUP TO
ENSURE THE FLOOR HAS THE LATEST RELEASE OF SOFTWARE.
2. Carrier pressure alarm configuration screen is not configurable for
primary and 2nd table to correspond with high and low end pressure
settings. Second table continues to error with low down force
settings of 80lbs with a 10% error limit. Need one screen for low
end and one for high end conditions.
WILL BE AVAILABLE IN SPEED495.
Process Issues:
1. System configured with old generic carriers/plates for process
testing. Why was the process portion not planned with new
customer carriers/plates?
SPEEDFAM FINDS IT VERY DIFFICULT TO USE CUSTOMERS' PLATES
BECAUSE THEY HAVE TO BE DIS-ASSEMBLED AND RE-ASSEMBLED PRIOR
TO SHIPPING WITH THE TOOL. TO CORRECT THIS ISSUE THEY WILL
HAVE AT LEAST 5 "NEW" CARRIERS ALWAYS BUILT AND READY FOR
CUSTOMER USE ON THE PRODUCTION FLOOR.
Process Results:
Parameter Spec Actual
============== ======= ========================
Avg removal rate >=2000 3729
Avg %Std.Dev. <=5.0 4.26
Max %Std.Dev <=8.0 8.1 (6.73 without 1st run)
Max HTHNU <=5.0 7.88 (4.59 without 1st run)
Throughput >35 wafers/hr 36
These results are consistent with those observed at previous source inspections
and the system is deemed acceptable pending the closure of the open issues
listed in this report. Speedfam has agreed, and assured us the open issue WILL
be closed prior to shipping. We will call to confirm on 12/18/96. The tool
will ship on 12/19 and arrive here on 12/23....Merry Christmass!
OTHER ISSUES AND ACCOMPLISHMENTS DURING THE VISIT AT SPEEDFAM:
o Signed-up Digital for Beta testing of Speedfam oxide endpoint system. We
will be the 3rd beta site and are scheduled to receive the system 3/1/97.
o Signed-up Digital for the 1st beta test of Speedfam W-Endpoint system.
Should come with the Auriga tool#65 in 2/97. Looked at their UEP system.
They've run into repeatability issues with this carrier current monitoring
system. Also heard some negative feedback on Luxitron system from AMD.
o Met with the engineering design manager to go over CESs for Auriga. These
included requirements and some design ideas for the 2nd platen pad
conditioner and slurry pump set-up.
o Met with Randy Smith, Software Engineer, we Demo the SECSII software package
and reviewed the VIRTUAL CONTROLLER design requirements. Both systems Look good,
we will need to work with CIM to get Digital's end of the interface software
working.
o Looked at their second platen wafer-loss detect system. Looks promising,
however, it is too big and needs to be miniaturized.
o Held a meeting with the Reliability Manager, the Filed Service manager,
Reliability Engineer, NE Field Service Manager, and the NE Sales Manager to
discus CMP.B1 head-3 stalling problem. Although Speedfam is working on a
long term rotation failure fix, I requested an immediate plan of action. It
was decide that an "old style" motor would be located, installed on head-3
and tested at 400lbs. The initial test failed, the motors's resolver
was than rotated 30 degrees and re-installed. Head-3 ran with out a
stall at 400lbs. A long term motor stalling plan is in progress at Phoenix,
due some time in January. This fix will apply to ALL CMP.B tools. It most
likely will include all new infranors, and some software constant changes.
o Meet with Bill Spencer, Carrier Films Manager, in Phoenix on Monday and
finalized the DIGITAL carrier list. There is only one plate mia. The updated
master list will be generated and put in carrier build room. We own and can
account for 61 plates, plus 20 new plates shipping with B4, for a total of
81 plates.
o Meet with Mike Slepekes, training manager, to discuss plan for Auriga
training. The first available class for customers is not until late March.
Once a schedule is finalized, it will be sent to Digital training dept.
|
23.3 | Install progress | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Mon Jan 13 1997 15:41 | 21 |
|
Install status for 1/13/97
KSI got alot of things done today. They are waiting
for an insert for the N2 line which should be done by this
afternoon. They are "drawing" up plans to construct a safety
barrier with unistrut to go around the plumbing. They are
waiting for a manifold set-up for in the fluids cabinet and
in the VMB to be assembled. They are planning on having
everything done for tomorrow's sign-off.
On our end... With SF Reps, the chiller cables were run
from the tool to the chiller and connected and the light
tower was re-mounted on top of the tool.
All of the facilities are ready at the valves on the wall
and we'll turn them on tomorrow to check for leaks.
The red CMP.B4 label was ordered for the front.
BJ
|
23.4 | vac pump in. | FABSIX::M_RIDLON | Eliminate the oblivious | Tue Jan 14 1997 06:23 | 4 |
| Vacuum pump is installed.
B-shift
|
23.5 | CMP.B4 install progress for 1/14/97 | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Tue Jan 14 1997 16:37 | 20 |
|
Install status for 1/14/97
KSI didn't have the parts in-house for the slurry manifolds
so we're going to have to wait another day on that. The did however
construct the unistrut frame to go around the plumbing but the
floor panels weren't ready so don't lean on it.
The chiller was filled with 50/50 glycol/DI. The system wasn't
powered up so we'll top it off when it does tomorrow. The PCW, DI,
N2, CDA, PV and RAW water were turned on. There was a tee fitting
that was used at the factory to flush the system of DI which we removed.
Rich and Dave installed the vent covers and other panels that need
to be installed after shipping.
We'll be firing up this puppy in the a.m.
BJ
|
23.6 | Install progress for 1/15 | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed Jan 15 1997 17:18 | 29 |
|
Install status for 1/15/97
Tool was powered up in the a.m. with one minor
issue. The Infranor for the conditioner suffered
electrical death during transit. That was replaced
with a new one.
The chiller was powered up and topped off. It's
now flowing fine throughout the tool lines.
The vacuum pump was pulled apart and cleaned out.
the reed valves were replaced.
Right now there is an issue with the monitor in front
not coming on. There is a issue with the SF A/B box inside
the tool. SF Reps are looking into it now. There is a
15 pin female to female adapter that is no where to be found
so we might have to do a road trip to Radio Shack tomorrow.
This connector is needed to connect the remote monitor to
the tool.
There is no door interlock installed on the door to the
electrical cabinet. This was an option for the customer. We
have asked SF to install one.
The off-shift transplant dude
|
23.7 | installation. | SUBPAC::CMP | | Thu Jan 16 1997 11:26 | 6 |
| Began installation of tool on Wed. found that the bleed valve for the
vacuum installed backwards, found a bad infranor for the pad cond.
(it was dead on power up) we are working on an issue on the switch
box (monitor at front of tool not responding), but at this time we
are able to use the remote monitor for the initialization. We will
continue with the installation and update.
|
23.8 | install continues | SUBPAC::CMP | | Thu Jan 16 1997 18:46 | 11 |
| Continued the installation today with these things completed.
checked and adjusted N2, air, and DI water pressure switches.
adjusted vacuum pressure switches on the carriers and vacuum tank.
checked and calibrated speeds and alignments for polish table,
scrubber, and carriers.
check and adjust epr regulators for the arms.
check carrier to cup alignments.
check cup heights.
We will be continuing tommorow and hope to get some wafers started.
|
23.9 | interlock installed. | FABSIX::L_SHKOLNIKOV | | Fri Jan 17 1997 05:39 | 9 |
|
This is just to let you know that electrical cabinet door has
an interlock now. It connected with the EMO button in the back
of the tool.
Leon.
|
23.10 | MECH ACCP PLAN 1/17/97 | STRATA::TDYER | | Fri Jan 17 1997 14:27 | 50 |
| +---+---+---+---+---+---+---+
| d | i | g | i | t | a | l |
+---+---+---+---+---+---+tm-+
(semiconductor)
I N T E R O F F I C E M E M O
To: FAB6 CMP Date: January 17th, 1997
cc: Dennis Goodwin From: Timothy P. Dyer
Dept: Equipment Engineering
Ext: 5072
Node: SCOMAN::TDyer
Lock/Mail Stop: HLO1-1/p-11
-----------------------------------------------------------------------------
CMP PASSDOWN
-----------------------------------------------------------------------------
**** I no longer have a long range pager ****
CMP.B1,2,3 - No issues. If head stalls occur on B1, document every possible
detail and give me a call
CMP.B4 - Mechanical acceptance testing will be run this weekend with a
barrels connected locally to each slurry channel. Once Speedfam
has completed their testing and set up, they will inform Brian J.
and he will start the marathon. The following test and recipes
will be run.
wafers recipe
5 bin-1
5 bin-2
5 bin-3
5 breakin
5 IO_MQC
5 IO_MQC
pre-read 5 12k ---> 5 IO_MQC post read and record, continue....
5 MECH_400 (400lbs)
5 MECH_400 (400lbs)
5 MECH_400 (400lbs)
450 MECH_30 (313lbs, 30 seconds)
------
TOTAL = 500
* TWO ASSIST ARE ALLOWED IN THE 500 WAFER MARATHON, IF A FAILURE
OCCURS, RE-START THE MARATHON AND CALL ME.
* PLEAS SEE BRIAN J. FOR MARATHON STATUS PASSDOWN, ONE DIGITAL
TECHNICIAN IS TO BE PRESENT DURING TEST.
(home phone 508 699-4905)
|
23.11 | install continues | SUBPAC::CMP | | Fri Jan 17 1997 20:36 | 6 |
| The continuing installation and power up saga.
We finished the alignments and checks, then started running wafers with
the no touch recipe. We managed to finish 35 before calling it quits
for the evening. We will start running tommorow with no touch recipe
to check operation then start on the marathon.
|
23.12 | Don't trash the carriers. | FABSIX::L_SHKOLNIKOV | | Sat Jan 18 1997 07:13 | 6 |
| I shut down the CDA supply for tonight to allow the carriers sat down
on the pad to keep them wet. Parts (video interface box and cables)
have arrived tonight and sitting on top of electrical cabinet.
Leon
|
23.13 | 500 WAFER MARATHON | STRATA::WESTPHAL | | Sat Jan 18 1997 19:54 | 72 |
|
CMP.B4 MARATHON
================================================================================
================================================================================
1st attempt
11:45 - 1 assist for scrubber not operating. Found a loose wire on break out
board #11. Decided to start the marathon over because we were only on
our 3rd run.
2nd attempt
15:30 - First assist for not picking up wafers from cassette B. Made
adjustments to the robot but suspect the problem was with the boat
itself.
16:45 - 2nd assist for not picking up wafers from cassette B. Reps. retaught
the robot and continued the marathon. Errors for cross slotted wafers.
CMP.B4's MQCs FROM RUN 7 2ND ATTEMPT
------------------------------------
particles
carrier large/total/scratch removal uniformity
#1 49 74 0 4032 12.1
#2 100 136 0 2716 19.9
#3 117 136 0 3078 16.2
#4 66 203 0 2760 12.3
#5 57 79 0 3219 15.0
----------------------------------------------------------
Removal mean 3161
Head-head range 1316
Uniformity mean 15.1
Uniformity range 7.8
CMP.B4 - Still having trouble picking wafers from cassette B. The SF Reps are
presently recalibrating the robot positions from scratch and will
than proceed with the 500 wafer marathon (starting at wafer #0)
Tim's note:
CMP.B4 - Mechanical acceptance testing will be run this weekend with a
barrels connected locally to each slurry channel. Once Speedfam
has completed their testing and set up, they will inform Brian J.
and he will start the marathon. The following test and recipes
will be run.
wafers recipe
5 bin-1
5 bin-2
5 bin-3
5 breakin
5 IO_MQC
5 IO_MQC
pre-read 5 12k ---> 5 IO_MQC post read and record, continue....
5 MECH_400 (400lbs)
5 MECH_400 (400lbs)
5 MECH_400 (400lbs)
450 MECH_30 (313lbs, 30 seconds)
------
TOTAL = 500
================================================================================
|
23.14 | B4 problems during Marathon | FABSIX::R_POIRIER | | Sun Jan 19 1997 07:07 | 16 |
|
Speedfam Reps started marathon and after 180 wafers
cassette "A" mapper errored for cross slotting. Reps
worked on problem for about 4 hours and problem went away.
Started marathon again and after 100 wafers, the same problem
occured. Reps determined problem is with the encoder for
cassette "A". They have ordered a new tilt assy. and expect it
in tonite (Sunday nite). Kevin will pick it up and bring it in.
Reps will not be back in until tonite (Sunday nite).
B shift
|
23.15 | Marathon run | FABSIX::R_POIRIER | | Mon Jan 20 1997 07:26 | 15 |
|
Speedfam reps replaced the tilt assmebly and tested
it. Worked fine. They had started the marathon and
had run 100 wafers when the Local evac sounded and
we had to leave the fab.
7:20 AM
We are back into the marathon. 150 wafers run, no errors
B Shift
|
23.16 | marathon continues | SUBPAC::CMP | | Mon Jan 20 1997 07:31 | 1 |
| Installed parts for the tilt unit, starting marathon again.
|
23.17 | UPDATE/PLAN | STRATA::TDYER | | Mon Jan 20 1997 13:53 | 27 |
|
CMP.B4 - We are currently 350 wafers into the 500 wafer marathon. the 1st
marathon was stopped at the ~200 mark for a bad encoder on the
A tilt module. Once the mechanical acceptance marathon is complete,
we we have several more items to address before releasing to Pro Eng.
1. Down force will be recalibrated based on the high head-head range,
seen in the data below. A repeat MQC will be performed.
2. Slurry manifold needs to be built and installed.
3. The Mega slurry system needs to be released in order to supply bulk
slurry to B4.
4. The SECS-II kit installation needs to be completed.
5. Local and global slurry detection needs to be installed and
tested.
carrier large/total/scratch removal uniformity
#1 49 74 0 4032 12.1
#2 100 136 0 2716 19.9
#3 117 136 0 3078 16.2
#4 66 203 0 2760 12.3
#5 57 79 0 3219 15.0
----------------------------------------------------------
Removal mean 3161
Head-head range 1316
Uniformity mean 15.1
Uniformity range 7.8
|
23.18 | Mechanical Acceptance Completed and Passed!!!! | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Mon Jan 20 1997 17:35 | 15 |
|
After the motor and encoder for A-tilt was installed
and tested, the marathon started. Everything went well.
The tool completed the 500 wafer marathon without a single
ASSIST. Rich is going to check the downforce on all (5) heads.
After they are checked, (5) MQC wafers will be run to check
the removal rate and uniformity. Jim Sadin will be running
the MQC's and will post the MQC data in CMP.B4 NOTES topic.
Leave the tool in WET MODE.
BJ
|
23.19 | CMP.B4 downforce cals | FABSIX::J_SADIN | Freedom isn't free. | Mon Jan 20 1997 18:33 | 18 |
|
CMP.B4 downforce cal info:
Head Before After
1 341.0 354.3
2 328.5 353.8
3 338.2 353.5
4 339.1 354.0
5 343.2 354.2
Pad Conditioner
before after
59.9 67.8
MQC data will be posted in the next note (currently running).
|
23.20 | marathon complete | SUBPAC::CMP | | Mon Jan 20 1997 19:11 | 4 |
| Fixed the tilt unit and started the 500 wafer marathon. The run
completed with no problems observed. Checked the down force again
and adjusted (point to point method), at this time MQC's are being
run to check tolerances.
|
23.21 | CMP.B4 mqc results | FABSIX::J_SADIN | Freedom isn't free. | Mon Jan 20 1997 19:53 | 16 |
|
MQC's completed on CMP.B4. Here are the results:
Head rem-rate %dev
1 3420 8.4
2 3430 11.9
3 3303 9.1
4 3368 7.9
5 3319 8.4
have a good night.
jim
|
23.22 | useless info | FABSIX::M_RIDLON | Eliminate the oblivious | Tue Jan 21 1997 01:09 | 8 |
| Tightened the locking bolts on the leveling feet (7 of 10 not secure)
and tightened 3 of the 10 feet so that they were making contact with
the foot pads. Also, whats up with the water level in the index table?
There's an 1/8" from the water to the top of the index table in the
front and in the back the water almost spills over the top.
B-shift
|
23.23 | B4 install update | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Tue Jan 21 1997 17:24 | 18 |
|
All of the panels were re-installed on the multihead.
The entire area was cleaned up.
The two almost empty barrels of slurry will be removed
from the area tomorrow. Two full ones might go as well.
We are still waiting for XT to make-up the manifold for
for the slurry in the fluids cabinet. They have been waiting
for parts since last week. After this gets installed, the
tool can be signed off.
We are hoping to get the mega system on line for next week
but nothing definite.
BJ
|
23.24 | CMP.B4 status.... | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed Jan 22 1997 19:05 | 22 |
|
The two empty slurry barrels were removed from the fab.
The two unopened barrels are going to stay for a while.
Process Eng. is contemplating using them for Process Acceptance.
Both sets of sliding doors were secured to the tool.
The parts for the slurry manifold are suppose to be in
tomorrow morning. If they are, KSI will have it on the tool
by tomorrow afternoon, Friday at the latest.
The front window panels were due in last week so they
should be here any day.
The stack of floor panels in the B3/B4 Maint. chase will be
moved under the floor for storage within the next week or so.
They all have cut-outs for the floor pedestals for B5 and B6.
They must know something I don't! ;^)
There is also a wall panel that will be removed shortly.
BJ
|
23.25 | | FABSIX::M_RIDLON | Eliminate the oblivious | Sun Jan 26 1997 06:56 | 6 |
| Installed Flouroware slurry detector module and ran associated cabling.
All thats left is to install the sensor once the manifold arrives.
B-shift
|
23.26 | Index table Infranor/moisture sensor | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed Feb 05 1997 06:00 | 11 |
|
While making the rounds, the tool was in WET MODE but
the is was alarming for INDEX TABLE not at home and the
moisture sensor was activated. The index table was traced
to a bad infranor and was replaced. All of the sensors were
checked but no sensors were tripped???
D shift
|
23.27 | Moisture sensor again | FABSIX::O_WHITFORD | Lead, follow or step aside. | Sat Feb 08 1997 03:59 | 18 |
| Same problem as last night. Found considerable puddle in the catch pan
under the unload elevator area. Lost tool power due to the door
interlocks and was unable to bring tool back up. Software is seemed
locked in DOS mode, would not respond to the keyboard, tried one last idea,
which got me to where I needed to be. Got tool initalized and we
are on the way to finding just where the leak is coming from.
Did find ONE drop hanging from one of the drain tube ports under the
scrubber. That was at the beginning of the investigation. Found a
significant leak around the left drain port in the scrubber bay.
Did some more looking and found that the leak appears to be under the
scrub wafer pedestal. A good amount of water was foound in the far end
of the transport paddle mount.
Dis-abled scrubber, removed brushes and put them in water track #1, put
tool back in wet idle.
D Shift
|
23.28 | Brushes dried up | FABSIX::B_JUBINVILLE | | Fri Feb 21 1997 06:44 | 9 |
|
Found DI water off on scrubber. Brushes were
replaced due the dehydration theory. They were
totally wasted!! Enabled the scrubber, got the
DI flowing and replaced all (6) brushes. All
set for process acceptance.
D shift
|
23.29 | N2 GUN | FABSIX::B_WESTPHAL | | Sat Feb 22 1997 19:58 | 5 |
|
CMP.B4 - Installed an N2 gun on the system.
|
23.30 | Slurry line, bracket repaired | FABSIX::B_WESTPHAL | | Sat Feb 22 1997 20:00 | 12 |
|
CMP.C1 - Kept blowing a slurry line off the SC112 filter housing. The line
that was blowing off was a thinner wall type tubing. Installed the
masterflex tubing we normally use and put a new 10 micron filter in.
Noticed that the slurry flow to table #2 was low but ran out of time.
the slurry flow still needs to be calibrated. The support that holds
the return pipe for the metal slurry broke off. Tried to polyweld it
several times using different rods but couldn't get a good weld. This
could be due to the different types of materials. Bolted the bracket
to the upper lip of the tank w/ ss screws. I will replace them later
when I find some Teflon screws.
|
23.31 | Hose broke on Slurry Pump #1 | FABSIX::R_POIRIER | | Tue Feb 25 1997 07:11 | 11 |
|
The hose on Slurry Pump #1 split open last nite while
in Wet Mode. DI water spilled onto the Moisture sensor
and the sub floor. Replaced the hose and put tool back
in Wet Mode.
B Shift
|
23.32 | "OOP's" I forgot! | SUBPAC::CMP | | Tue Feb 25 1997 08:47 | 2 |
| Replaced the slurry pump tubes on B4on 24 Feb 97. The tubes were worn.
|
23.33 | conditioner | FABSIX::R_GEE | | Tue Feb 25 1997 18:46 | 7 |
| Errored for "conditioner not arriving at polish table". Adjusted
the sensor up.
Errored for "final polish pressure out of tolerance". Set the
pressure ramp in segment #2 of the polytex final polish program
to 2 seconds.
A shift
|
23.34 | loose ribbon cable | FABSIX::B_FINN | | Mon Mar 03 1997 19:34 | 18 |
|
System faulted for no slurry. Found large air bubbles in the
line. Bob called Ashland for the mega slurry system. They had just
added slurry. Ran slurry in manual to remove air bubbles, aok since.
System errored for vacuum. Found double load on all five heads.
Ran again and watched, found that wafers were not being blown off the
carriers prior to loading the next 5 wafers. We ran and monitored i/o and
found vacuum turning off prior to blow off but the main n2 blow off
solenoid (i/o 14) for N2 was not being told to turn on. Check fuse and
opto 22 aok. Problem was loose ribbon cable. Tested aok.
A shift and Speedfam
|
23.35 | regulator problem | SUBPAC::CMP | | Wed Mar 05 1997 18:49 | 3 |
| Worked with a problem with uniformity and found a regulator to be
intermitent. Ordered one from stock and found nothing there. Ordered
2 for tomorrow morning delivery.
|
23.36 | carrier one e.p. not responding | FABSIX::J_SWENSON | | Wed Mar 05 1997 18:52 | 9 |
| Carrier one E.P.
--------------
Carrier one E.P. not responding.
One has been ordered, tool down awaiting part's.
c-shift
|
23.37 | removal rate fix | SUBPAC::CMP | | Fri Mar 07 1997 16:23 | 4 |
| Operator noticed the removal rate was off slightly but within spec.
Replaced the pad and did down force cal for pad conditioner. Also
discovered the chiller was off. Complete the work and return to
process eng.
|
23.38 | A and B tilts adjusted | FABSIX::J_SWENSON | | Sat Mar 08 1997 19:21 | 24 |
| Tilt A and Tilt B problems
--------------------------
- Cassette A, Robot detecting all wafers to be cross slotted.
found front to back tension assy to be loose, adjusted
accordingly and ran 100 wafers without incident.
- Cassette B, Robot not picking up the 15th wafer and hitting
every wafer there after.
Adjusted the tilt front to back. Ran 100 verification wafers
with no problem.
*** Need to keep an eye on Cassette B, It seemed very strange
that it would not pick up only the 15th wafer and every
wafer prior was o.k. and every wafer after the end efector
would just nick the wafers. Tried several different
cassettes and the results were the same.
C-Shift
|
23.39 | regulator replaced | SUBPAC::CMP | | Tue Mar 11 1997 15:37 | 2 |
| Install EP regulator and adjust down force on all carriers. Return to
operation.
|
23.40 | Tilt Adj. loc tighted | FABSIX::J_SWENSON | | Thu Mar 13 1997 19:42 | 11 |
|
Cross slotted wafers detected @ A cassette
------------------------------------------
- Adjusted tilt and loc tighted adj. screw
- ran verification wafers without any problems
tool ran without any cross slotting probems.
C-Shift/Dave, Kevin
|
23.41 | Slurry pump 3 | FABSIX::J_SWENSON | | Thu Mar 20 1997 17:19 | 51 |
|
SLURRY PUMP 3
-------------
Slurry pump 3 has been hooked up temporarily without the presure
guage on the line.
The guage and flared tee are in the speedfam cabinet.
The valve off the manifold for slurry pump 3 has been turned off so
before using pump 3 the valve will need to be turned back on.
C-Shift
|
23.42 | VARIOUS ERRORS | FABSIX::B_WESTPHAL | | Thu Mar 20 1997 19:51 | 9 |
|
CMP.B4 - Never went down in the system but the tool was coming up with strange
errors. When production tried to run some tests for engineering they
got various errors. Soft booted the tool and while it was homing I
received a vacuum error. The gauges on either side of the filter had
different readings so I changed the vacuum filter. The gauges still
were still reading 4lbs. Different so I don't trust the gauges.
Hard booted the tool without any errors and ran 5 wafers on the
IO_MQC recipe without any trouble. Logged an assist.
|
23.43 | DI dripping under scrubber | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Sat Mar 22 1997 05:59 | 14 |
|
The moisture alarm was going off after lunch. Traced
problem to the sensor under the scrubber area. Water is
dripping from the same area as just after the install.
The water is wiped up but we will monitor the leak. The
area will most likely need RTV to seal it up. I will be
forwarding this note the SF dudes as this is the second time
for this.
D shift
|
23.44 | Automation frame di water leak / slurry pump #2 | FABSIX::R_GEE | | Mon Mar 24 1997 08:54 | 13 |
| There was a DI water leak coming from under the scrub station.
Found that the leak was coming from the scrub station wafer
pedestal area.
Removed the scrub station wafer pedestal assembly, Teflon taped
the DI water fittings, cleaned the o-rings, silicon sealed the bolt
feedthru's and re-installed the assemble. No more leak.
============================================
Slurry pump #2 hose ruptured, replaced it.
A shift
|
23.45 | DI hose rupture | FABSIX::J_SWENSON | | Sat Mar 29 1997 19:19 | 27 |
| Di line rupture
---------------
- Found di line ruptured in pump box.
- Replaced all three lines
- moisture sensor was under water and remained green.
- will pass info of the moisture sensor to B-Shift.
NOTE:
- noticed vacuum line to bulkhead fitting to be the thin wall type
which we had problems of colapsing on other B tools.
- noticed DI line which feeds through the flex track to be the same
type which ruptured on B2.
C-Shift
|
23.47 | installation | SUBPAC::CMP | | Thu Apr 03 1997 21:21 | 3 |
| Started the installation of the tool this week, all was going smoothly
until this afternoon when the GFIC relay tripped causing the tool to
shut down. Parts are on order for tomorrow when we will continue.
|
23.47 | guage installed | FABSIX::J_SWENSON | | Sat Apr 05 1997 19:52 | 13 |
| New guage
----------
Installed new guage, guage guard setup
on line 3.
Pressure when not running slurry
0-15 psi
C-Shift
|
23.48 | Mapper Errors / Wafer Breakage | ASDG::POIRIER | | Sat Apr 12 1997 20:43 | 66 |
|
Mapper Error:
------------
Tilt Mechanisim #1 repeatedly created mapper errors this morning.
We sucessfully mapped a full cassette of wafers with no problem.
We then sucessfully mapped using 5 different full cassettes.
Changing cassettes didn't seem to make a difference. The
mapper error reoccured everytime there were 15 wafers in a
cassette. It wouldn't with more than 20 wafers, or less than 10.
It repeatedly failed with 10 to 20 wafers in ANY cassette.
We noticed that the 3 spring loaded leveling screws under the tilt
mechanisim platform looked different. The springs on the 2 screws
facing the front of the automation frame were fully compressed
while the spring on the 1 screw that faces the rear of the automation
frame was barely compressed. This condition requires the weight of 10 or
more wafers to compress the rear spring to the point where the
Robot will sucessfully map the cassette.
We're theorizing that the tilt mechanisim platform needs to be readjusted
to a level position while maintaining equal compression on the 3 spring
loaded leveling screws.
We didn't want to change the tilt mechanisim adjustment since none
of us knew how to teach the new style robot new posistions.
We managed to get through the shift by asking gwen to load more than
10 wafers into a cassette on each run. She then selected to polish
wafers from the appropriate slots.
Crash at 7:00 P.M. :
------------------
The system broke 2 product wafers at 7:00. We're not sure when the
breakage occured because there were no error messages or audible
alarms.
This was the condition the tool was in when the breakage
was first noticed by operations:
1) Head #4 down on table, all other heads up.
2) Wafer missing out of head #4. Wafer broken on pad with chips
under head #3.
3) Wafer in head #3 scratched and cracked in half but still in
carrier.
4) Blue stripe across bottom of touch screen with words " ENDING RUN".
5) The tool was running the IO_ILD1 recipe at the time of the
crash. The tool was 10 seconds into segment #7. Gwen informed
us that we were half was thru the 20 sec rinse at 313lbs of
down force and that she had been running this recipe all day
no problems.
We're currenly cleaning up the mess and removing the carriers
from the tool. We'll be handing over the troubleshooting and
verification to "B" Shift shortly.
P.S Thanks gwen for assisting us in the data collection.
C Shift.
|
23.49 | Load station problems | FABSIX::R_POIRIER | | Sun Apr 13 1997 08:48 | 17 |
|
We had some problems with the load station.
Load 1 kept erroring for cross-slotted wafers.
We leveled the load platform and were able to
clear this problem.
Then a problem occured with Load 2. The robot
is going in to far to get the wafer and putting
it the pre-aligner off center. The pre-aligner
doesn't know it has a wafer and when the fingers
come down to get the wafer it can't pick it up.
Didn't have the time to make necessary adjustments.
B Shift
|
23.50 | CMP.B4 | FABSIX::J_SADIN | Freedom isn't free. | Sun Apr 13 1997 20:42 | 17 |
|
Had wafer crash out of head 2. Cleaned up machine, redid
spindle/cup alignments (2 and 3 need copious adjustment) and placed new
primary pad and one new carrier (#2) on system. Found 2 of 3 drive
belts for main table are loose/sloppy (reps will need to address this
issue). Left as is (no adjustments made)..table performance was ok.
Unload flipper was driving past wafer present position (closed).
Adjusted fingers..working ok now. Door interlock on unload side front
is not working (bimba will not drive piston properly...binding up
inside). Bimba seems to be ok, won't work when installed in tool.
Passed on to b-shift.
a_shift
|
23.51 | fixes and adjustments | SUBPAC::CMP | | Mon Apr 14 1997 13:43 | 8 |
| Checked into some complaints about operation and found that the bimba
on the door was OK but the lock shaft was bent, replaced with a new
one, old one was straightened and should be good for the next time.
Polish table seems to have excess play, found the belts were loose,
retightened pulley assy. belts are now tensioned good. Robot had mapper
errors, we found the adjustment screws were loose, put locktight and
readjusted(screws were ones that did not have loctight). right now
tool is in wet mode waiting for process work.
|
23.52 | crash, program edited | SUBPAC::LANDRY | | Mon Apr 14 1997 20:46 | 30 |
| Well, another rough ride today with this system......
Several nuances from the previous day were taken care of early
in the day. See the previous note for details. After all items
were fixed, we began to perform the breakin of the pad that never
quite got done the previous day. All the respective "BREAK IN"
recipes ran flawlessly. We turned the system over to Opts for
MQC's. The very first run, the system crashed. A wafer got away
from Head #1, taking carrier #5 all with it. Cleaned up the mess.
RE-checked head to cup alignment, which we confirmed as good. Down
force was check on all heads. All were within 1 pound of spec except
for head #1, which was 3 pounds out. Regardless, all were tweaked
to tighten things up a bit (seeing we were there anyway).
After all the above checks were completed, we scrutinized the IO_MQC
program for discrepencies. We found that the post polish rinse steps
showed that the program was delivering 40 second on rinse water with
300+ pounds down force. The program was adjusted (segment 7 was
changed) for the down force to change in segment 7 from 313 pound
down force, to 80 pounds down force. Therefore, segment 6 is 20
seconds DI at 377 pounds, and segment 7 at 20 second DI at 80 pounds
down force.
MQC's to be run soon. Unfortunately, a contractor messed up on some
plumbing lines, which caused the slurry in the tank to get diluted.
We are currently in hold (ON B4 ONLY, AS IT'S THE ONLY SYSTEM UTILIZING
THIS SLURRY), awaiting word from Ashland Chemical that the system is
ready to run.
A shift
|
23.53 | adjust prealigner | SUBPAC::CMP | | Tue Apr 15 1997 20:10 | 3 |
| The operators noticed the wafer being hit with the load flipper.
When check on found the prealigner did not align wafer. Made an
adjustment to the air valve to move fingers.
|
23.54 | DOUBLE LOADED A WAFER/SCRATCHES | FABSIX::B_WESTPHAL | | Thu Apr 17 1997 20:30 | 11 |
|
CMP.B4- Down for "carriers not rotating" error, the screen being locked up and
scratches. There were 2 wafers on carrier #1 and wafer #1 was missing
from unload cup #1. Obviously carrier # 1 failed to blow the wafer off
from the previous run. We had the same problem on CMP.B1 last week.
Recovered wafers, rebooted the tool and observed tool while production
ran 10 warm ups and 5 MQC wafers. Tool went down later for scratches.
Cleaned the conditioner and process area, changed both pads and
carriers #1 & #2. Dumped the index table, pulled the cups and wiped it
out with DI wipes. Passed all MQCs.
|
23.55 | PROBLEMS W/ CARRIER #2 | FABSIX::B_WESTPHAL | | Fri Apr 18 1997 21:49 | 10 |
|
CMP.B4- The tool never went down in the system but it was not blowing the wafer
off carrier #2 very well. It was noticed that Air was bubbling out
between the secondary housing and the gimbal assembly. Took the
carrier apart and found that the carrier plate was not tight against
the secondary housing even though the bolts were torqued to 8 lbs.
With process engineering's permission the bolts were tweaked a little
to pull the plate against the housing. This corrected the condition
and the tool was returned to engineering.
|
23.56 | IF PRODUCTION NEEDS CMP.B4 | FABSIX::B_WESTPHAL | | Fri Apr 18 1997 21:50 | 10 |
|
* Note Concerning CMP.B4 from Bruce Tower- CMP.B1 was put to external to begin
decontamination and preparation for shipment back to SPEEDFAM. Engineering
would like to keep CMP.B4 as long as the other tools are running okay. In the
event that production should need the tool Bruce requested that slurry pumps 1
& 2 be hooked up to a barrel of SC112 (which is at the tool). The proper
tubing to hook the barrel up to slurry pump #1 & #2 solenoids are on top of the
barrel. This is because the MEGA SLURRY SYSTEM's percent solids is at 18%.
An MQC must be done to verify there is no problems w/ the slurry.
|
23.57 | Need SC112 drum hooked up for Ops. | STRATA::BTOWER | | Sat Apr 19 1997 12:45 | 10 |
| The last 2 test lots using SC1 blended to 18% should be done today.
Gwen will let Brad know for sure when that happens. We will not be
Testing SS25 at 18% until Tuesday. So, we need to hook the SC112
drum up to the system and turn over to production for the rest of the
weekend. Please keep a close eye on the barrell to make sure there is
enough before Ops starts each lot. The system was fully qualified
at 12.8% solids before we started the other slurry tests. Process Eng
has also sent out a note stating such. Of course the Ops folks will
run MQC's again to make sure of system integrity before any product is
run through it. Any questions or concerns give me a call at home.
|
23.58 | Tool now running on 12% slurry from drum. | ASDG::POIRIER | | Sat Apr 19 1997 20:50 | 24 |
|
Slurry Drum:
-----------
The tool is now running on 12% slurry from a 55 gal drum in the
service chase. Keep an eye on the level!
Carrier #2 Blow Off Problem:
---------------------------
Carrier #2 was changed due to the wafer not blowing off. The carrier
was replaced and no further problems with the blow off were observed.
The carrier was returned to Sammy for tear down and inspection.
C Shift.
Slurry build up on Carriers:
----------------------------
Carriers #4 & #5 had to be scrubbed again today due to slurry build up.
We've been cleaning #4 & #5 daily for this condition.
C Shift.
|
23.59 | Failed MQC's for particles | FABSIX::R_POIRIER | | Sun Apr 20 1997 08:40 | 14 |
|
B4 failed MQC's for high particles. We replaced
the pad. Also found that the slurry pump #1 had
blown. The moisture sensor had not alarmed because
a rag was founf under the sensor used instead of
the proper pad. We couldn't find any pads in the
room. SGMR was down. Unable to order more pads.
Replaced tubing for pump #1 and put another rag
under the sensor.
B Shift
|
23.60 | CMP.B4 passdown | FABSIX::J_SADIN | Freedom isn't free. | Sun Apr 20 1997 20:37 | 22 |
|
Machine was vibrating during a run and it was noticed that no
slurry was flowing out of the dispersion tubes. Checked drum and found
it to be full. Checked the slurry pumps and found the lines on all
three peristaltic pumps were cracked and leaking profusely (leak
detector didn't alarm?). Replaced all the lines and placed a piece of
"leak detector paper" under the leak detector. Slurry flow was
restored.
Had another error for "scrubber not in". Checked sensor and it was
being met. Ran a few test runs without another error. Not sure what is
up with that.
Hooked up slurry loss detector to the slurry lines being used out
of the barrel....it works, too (we ran out of slurry late in the day and
it set off the sensor). Replaced drum with a fresh one.
a-shift
|
23.61 | door sw | FABSIX::B_FINN | | Mon Apr 21 1997 20:47 | 17 |
|
System had vibration during polish. Operator EMO'ed the tool. Doc and I
were downstairs at ashlands. No slurry or other alarms. We started up tool.
It would not initilize due to open process doors. All doors were shut. Found
intermittent double door sw. Removed and repaired. aok.
Brought up new slurry barrel. Installed at 5pm and brought empties back for
5 cent deposit.
A shift and Speedfam
|
23.62 | slurry barrel | SUBPAC::LANDRY | | Tue Apr 22 1997 20:31 | 5 |
| Put a new barrel of slurry onto the system today6 at 1:30 PM. The Mega
system is still not ready to be turned on at this point.
A shift
|
23.63 | Froze-Up at segment 1 twice | FABSIX::A_QUILLEN | Nobody will get away with anything | Wed Apr 23 1997 08:11 | 8 |
| During the shift, system went down twice. It froze-up at segment 1 with
the screen indicating "wating for carrier 1 to home". The first time
system was soft booted and tested fine. We greeted second incident
with an EMO. Checked voltages on the infanors. Re-seated cables to the
CPU and DCX cards. Re-booted and ran 40 test wafers without a problem.
Back to production.
B shift
|
23.64 | waf detect/carrier position | FABSIX::B_FINN | | Wed Apr 23 1997 20:39 | 26 |
|
System continued intermittent errors first notice on b shift, where system
would fault for carrier not at home as soon as the multi head reached the
polish table with wafers (always second run of wafers on touchup b).
No hardware errors found, whilst trying other recipes and deleting and
rewriting touchup b, error changed to "time out waiting for carriers at clean
station". Keep in mind this is seconds into segment 1. Traced k2 power fail
back to wafer loss board. When we jumpered out wafer loss, no errors occurred.
Frequency of error increased to every run and error was finally captured by
wafer loss board and IO35 as a momentary intermittent wafer loss detect signal.
adjusted sensor 5.
We are now repeating I/O MQC and touch up recipes to verify the fix since
the problem changed and evolved into a wafer detect error.
System is still running off the local slurry barrel, not the house system.
New barrel of slurry was installed today. You will need to monitor its
level during the night.
A shift and Speedfam
|
23.65 | Conditioner, pads, infranors oh my... | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Thu Apr 24 1997 04:56 | 20 |
|
Christened the tool for it's first PM. I'm going to
start inverting money in the Infranor business. We had
the tool powered down for 20 minutes or so and we were
looking into a leak in the drip pan under the scrub station.
When we powered the tool back up, we had L2 lit up on
the Infranor for head #3 rotation. That was replaced and
things were running again. When we were taking the conditioner
of the tool, 4 out 6 screws/helicoils were completely out of
conditioning disk adapter. How this didn't fall off is beyond
me. Anyways, this tool has the new teflon/plastic adapter ring.
The only ones in stock were the metal style. We installed the
metal style ring adapter and had to re-cal the conditioner
downforce because of the weight difference. After we brought
the tool up, it was discovered there were over 2000 wafers on
both pads. Both pads were replaced and the tool was turned
back over to OPS.
D shift
|
23.65 | Infranors, pads and flippers oh my...... | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Thu Apr 24 1997 08:41 | 27 |
| Christened the tool for it's first PM. I'm going to
start inverting money in the Infranor business. We had
the tool powered down for 20 minutes or so and we were
looking into a leak in the drip pan under the scrub station.
When we powered the tool back up, we had L2 lit up on
the Infranor for head #3 rotation. That was replaced and
things were running again. When we were taking the conditioner
of the tool, 4 out 6 screws/helicoils were completely out of
conditioning disk adapter. How this didn't fall off is beyond
me. Anyways, this tool has the new teflon/plastic adapter ring.
The only ones in stock were the metal style. We installed the
metal style ring adapter and had to re-cal the conditioner
downforce because of the weight difference. After we brought
the tool up, it was discovered there were over 2000 wafers on
both pads. Both pads were replaced and the tool was turned
back over to OPS. Unfortunately, the tool never got off the
ground. The unload flipper had a wafer but kept alarming because
the sensor wasn't being made. We calibrated the unload flipper arm
and seems to be ok now.
on a side note... the robot is either going into get the wafer
too far or it doesn't go in enough at the pre-aligner. It will
need to be cal'ed. The pre-aligner is centering it ok though.
D shift
|
23.66 | NO SLURRY | FABSIX::B_WESTPHAL | | Thu Apr 24 1997 21:08 | 10 |
|
CMP.B4 - An assist was made for a "no slurry" error. The tubing floated up in
the barrel so it could not draw slurry. This I believe damaged the
wafers that were polishing. Tied some more weight on the line,
cleared the alarms soft booted the tool and returned it to production.
|
23.67 | Mega slurry system repaired. New Tank on order. | LUDWIG::BTOWER | | Fri Apr 25 1997 16:42 | 10 |
| The Mega slurry tank has been welded and tested for leaks. Looks good
right now. Ashland Chemical is currently in the process of flushing
Loop 1(B4) and Loop 2(Auriga) of water. In about an hour they should
have slurry for us. It may be slightly off the normal 12.8% but close.
They will be beeping John Swenson when they are all set. Please remove
the SC112 barrel and hook us back up to the Mega system. OPS should
run MQC's again.
FYI: Monday we will begin testing of SS25 diluted to 18% out of a drum
on this system.
|
23.68 | POOR UNIFORMITY/SLURRY NEEDS TO BE HOOKED UP | FABSIX::B_WESTPHAL | | Fri Apr 25 1997 21:00 | 11 |
|
CMP.B4 - Down for poor uniformity on all heads. The uniformities were in the
20s so we changed the pad and put the tool to verify. The tool came
back to us for the same thing. The pad conditioner down force and
centering. The down force was at 78 so that was left alone and the
centering looked pretty good but that was moved from 9.2 (ID & OD) to
9.1 (ID & OD). A new primary pad was put on for good measure. The
tool sets needing conditioning, breakin, and MQCs. The okay was also
given to hook back up to the MEGA slurry system so that also needs to
be done.
|
23.69 | Nothing but trouble. | FABSIX::L_SHKOLNIKOV | | Sat Apr 26 1997 08:49 | 15 |
| The tool was hooked up to the mega system and trouble has begun.
First of all we have tried to calibrate the slurry flow and we couldn't
get it right. Found inconsistent slurry pressure out of mega system.
Ashland has been called and problem corrected for a while. The slurry
pumps finally were calibrated at 550 ml each. The slurry detection
system was verified after that and it did not responded at all. The
sensor seems to be worked but the control unit not, so this one was
replaced with a new one. The tool was released to manufacturing and
passed mqc's with removal rate ~3300 and st.dev. 7-8%.First product
send ahead 5 wafers have been run with r/r 4600 st.dev 12-18% and
dead center . At this point we decided to replaced all 5 carriers and
check it again. All 5 carriers which were removed from the tool got
guide rings polished all way down.
d - shift.
|
23.70 | AIR IN MEGA SYSTEM CAUSES SLURRY TO SPUTTER | LUDWIG::STANDING | | Sat Apr 26 1997 20:16 | 6 |
| CMP.B4 Went down for not distubuing slurry evenly, air in line (MEGA SLURRY
SYSTEM). Ashland was called. They found both filters were online.
One was taken off line. You can still see an air bubble in the
slurry line right before SV70,71. Ran pumps to purge out any air
that still might be entrapped in the line. Changed pad due to poor
uniformity, and changed carriers 3,4 because poor uniformity.
|
23.71 | vibration | FABSIX::B_FINN | | Mon Apr 28 1997 20:14 | 24 |
|
Vibration occurred on Ild3 recipe. Ops tech EMO'ed the tool, segment and
time was not caught prior to EMO. We were paged from the metals room.
Started tool up and unloaded product. No breakage. Slurry was present at input
to slurry valves and Mega system computer showed no recent history of alarms.
Ran slurry in manual aok. Ran notouch recipe and monitored i/o's and valves for
proper operation. Found pump 1 shutting off intermittently in recipe then
restarting on its own. Tried different recipes still errored. Edited Notouch
recipe to toggle pumps on and off. Swapped pump 1 controls at molex connection.
problem swapped to pump two. Set pump molex connections back to original
positions, and swapped i/o inputs to pump controllers. Problem swapped to pump
two. Changed Analog output board 7. Rich also lowered all pump setting below
1000. Ran 4 runs including ILD3 recipe aok. Need to monitor due to its
intermittent nature.
A shift and Rich
|
23.72 | ACTION ITEMS FOR SPEEDFAM | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Wed Apr 30 1997 07:35 | 12 |
|
Some warranty items for Speedfam....
1) continuous leak under scrubber into drip pan
2) robot to pre-aligner needs to be adjusted
on a side note - in HARDWARE.CMP, Check Main Vacuum Line
was set to 0. If we lost main vacuum, it would never alarm
for it.
|
23.73 | slurry alarm. | FABSIX::L_SHKOLNIKOV | | Thu May 01 1997 06:44 | 5 |
| The tool was alarming for no slurry available, found air bubbles in
the line. Run pumps to purge out any air that still could be in the
line. No problems since then.
D - shift.
|
23.74 | musical slurry | LUDWIG::BTOWER | | Mon May 05 1997 17:24 | 5 |
| Matt should be able to complete the SS25 (18%) test lot today. When
his lot is complete we need to remove the barrel and hook the system
back up to the mega system. Please keep the barrel of SS25 in the room.
The lot that we ran Saturday will be back for ILD3 later this week or
early next and we'll need to hook the ss25 back up.
|
23.75 | re_seal as best we can please | LUDWIG::BTOWER | | Mon May 05 1997 17:30 | 3 |
| Addendum to previous note: Please re_cap or tie wrap a glove or
something to the opening in the barrel to keep air out as best we
can. Thanks....................BT
|
23.76 | Wafer loss sensor adjusted (head #5) | ASDG::POIRIER | | Thu May 08 1997 20:21 | 8 |
|
Wafer loss error:
-----------------
The system errored for wafer loss on head #5 while in seg #1 of
the t_UP_A recipe. The sensitivity on head #5 wafer loss sensor was
adjusted. There were no futher problems or the rest of the shift.
C Shift.
|
23.77 | Wafer detect sensor loss | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri May 09 1997 06:18 | 8 |
|
Tool failed for wafer detect system. Found #3 and #5
wafer loss sensors tripped on the board. The sensors
were not hitting anything, they just needed re-calibration.
D shift
|
23.78 | Elevator#2. | FABSIX::L_SHKOLNIKOV | | Fri May 09 1997 07:27 | 5 |
| Tool went down for crossloting wafers in elevator #2.
We adjust the water level in the elevator and run 25 wafers
to verify. Tool returned to production.
D - shift.
|
23.79 | Unload flipper box sealed & adjusted / Srubber station leak / high particles | ASDG::POIRIER | | Fri May 09 1997 20:29 | 38 |
|
Unload Flipper to Unload Cup Alignment:
--------------------------------------
The unload flipper box cover was removed and sealed with rtv.
After the box was closed up, the flipper fingers were moved which
caused a wafer pick-up misalignment. The flipper to cup and flipper
to scrub aligments were reset. A 15 wafer verification run was
completed with no further problems.
Scrub Station leak:
------------------
The leak under the scrub station was sealed with RTV this afternoon.
The RTV cured for 2 hours before the water was turned back on.
We just checked it and its now leaking about 2 inches away from the
sealed area. It looks like another application of RTV will do the job.
Put a bead of RTV around the cut out under the scrubber area.
(A green air line is located in the center of the cut out)
High Particles:
--------------
The system went down at 6:30 P.M. for Pad step height and HIGH
PARTICLES.
1) We noticed a small quantity of rust particles around the
inside edge of the pad. A closer inspection revealed rust comming out
of the center of the platen. (At the junction of center ring and the
main platen) We cleaned with a scotch pad and rinsed with DI water.
2) The Di water pressure was at the low end of spec. 20 psi
The spec is 20 -30. It was increased to 29 psi.
The gage on the wall is reads zero.
B2 = 27 psi
B3 = 30 psi
|
23.80 | Load flipper errors | ASDG::MORIN | The Quest is on | Sat May 10 1997 06:39 | 7 |
| System went down for wafer present sensor of load flipper not being
made. Checked the digital i/o's,lights intermittent. Opened upper
assembley of the flipper box and adjusted the sensor block and
lubricated the mechanical parts. Ran 25 wafers for verify and returned
back to production.
D-shift
|
23.81 | Input 70 | FABSIX::J_SWENSON | | Sat May 10 1997 19:38 | 15 |
| All Clamps Not Up
-----------------
- found input 70 not cooperating.
- clamp 2 upper switch not making.
- adj. switch
- 20 or so multihead moves no problem found.
C-Shift
|
23.82 | safety upgrade | SUBPAC::CMP | | Tue May 13 1997 20:25 | 3 |
| Today we installed a safety lock for the multi-head. To use this
device, you place it under the multi-head lock lever when in the locked
position. This prevents the multi-head from moving.
|
23.83 | carrier 3/robot problem. | FABSIX::J_SWENSON | | Thu May 15 1997 20:34 | 16 |
| carrier three slow on coming up.
--------------------------------
Trouble shot 24vdc solenoid while disconecting and connecting
lines and testing the cda loop problem disapeared and tool was
returned.
Robot missing wafers intermittently on cassette B.
Tightened tilt plate screw and retaught robot pos for cassette B
and also noticed that the robot was not placeing the wafer correctly
at the alighn station so retaught that position.
Everything tested fine so returned tool back to production.
C-shift and Rich
|
23.84 | Rotation tolerance error | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri May 16 1997 22:40 | 7 |
|
Tool alarmed for rotation tolerance out of spec. Inspected
rotation on arms and table. No problems observed. Cycled
power to the tool to clear out any software bugs and 25 wafers
for VERIFY. Good to go...
D shift
|
23.85 | HIGH PARTICLES | FABSIX::B_WESTPHAL | Bradley BP#661 EXT# 6610 | Sat May 17 1997 20:52 | 5 |
|
CMP.B4 - Down for high particles. Particles were over 1300 for each head.
Ashland chemicals was contacted and a filter change was done.
Particles came in under 300 per head.
|
23.86 | High particles | FABSIX::R_GEE | | Sun May 18 1997 20:18 | 6 |
| The tool failed MQC's for high particles. We suspect the filter
for the mega system needs to be changed, this took care of the
particle issue yesterday. Called 3 different Ashland beeper #'s and
left a message at 6112, with no response back. Will have "B"
shift follow up with Ashland.
A shift
|
23.87 | scratches | FABSIX::J_SADIN | Freedom isn't free. | Tue May 20 1997 13:56 | 12 |
|
found severe scratching on five product wafers run on CMP.B4. Tried
to isolate the cause by running wafers only on the primary table, and
then only on the secondary table. Neither produced scratches. Cleaned
up all the carriers and the conditioner. Replaced the scrub station
brushes as they looked a bit dirty/worn. Ran a set of mqc wafers and
particles came out great (in spec, no scratches). Turned tool back over
to production.
a-shift
|
23.88 | Slurry barrell replaced | FABSIX::R_POIRIER | | Wed May 21 1997 07:30 | 8 |
| Replaced Slurry barrell at 6:15 am.
Ran slurry in manual to get rid of
any air in the lines. Gave tool back
to OPs.
B shift
|
23.89 | Barrel swap | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Thu May 22 1997 06:01 | 7 |
|
Drum on tool almost out. We put a new barrel
on and purged the lines. MQC's are due around 9am
so C shift, please put the tool back on the Mega-system
and check flow at that time.
D shift
|
23.90 | Slurry barrel swap-out | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri May 23 1997 05:59 | 8 |
|
Slurry barrel was real low and OPS was throwing
on a metal-3 lot so the barrel was replaced.
D shift
|
23.91 | Slurry barrell | FABSIX::R_POIRIER | | Sun May 25 1997 00:55 | 5 |
| Slurry running low. Replaced barrell at 9 PM.
B shift
|
23.92 | Infranor, PM and unload flipper issue | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Wed May 28 1997 08:28 | 33 |
|
Slurry barrel was getting low so it was replaced.
During the run, it appeared that the unload flipper
had the 3rd wafer of the run and some how dropped it
in the paddle track. The tool was alarming when the
operator noticed that the paddle was trying to go home
with the product wafer in its way and the tool was alarming
for wafer not present at unload. The tool was EMO'ed
while it was still polishing. All of the load and unload
wafers were removed and the tool powered back-up. Tool
alarmed for heads #2, #4 and #5 not rotating. All (3)
infranors had L2 lit on them. Multihead transport was
also red with L2 lit up. After another reboot, only
#4 wasn't rotating. The infranor was replaced and the
tool powered back up. Now only @2, #5 and MHTA where lit
with L2. After another powerdown, just #5 rotation was
lit with L2. There were no more in stock. Rich was called
and is trying to get (3) in here ASAP. OPS gave us the
tool for the week-late monthly. All of the PM's were
completed for the week on the tool. Head #5 was taken
off line and the carrier in staying wet. A WARNING will
will be placed on tool. The leak was investigated but still
no source found.
The unload flipper sensor block was loose and needed to
be tightened. Wafers were run for VERIFY. Tool ready for
break-ins.
D shift
|
23.93 | Particles and awaiting infranor for head 5 | LUDWIG::BTOWER | | Wed May 28 1997 18:40 | 9 |
| Spoke with Ashland this morning. They felt the mega slurry system
should be clean after multiple days of flushing it out. Removed barrel
and ran MQC's hooked to mega system. Particles still high. "Large"
around 300 and total 1100 on 3 out of 4 heads. Hooked barrel back up but
have had problems getting an MQC done due to scrubber issues. Gwen
should be doing one now. The infranors were ordered and should be here
"around" 11:30 tonight. Plan is to run 4 heads until the infranor's
show up.
|
23.94 | Barrel change | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri May 30 1997 07:58 | 4 |
|
Replaced slurry supply barrel at 9pm
D shift
|
23.95 | heater off on chiller | FABSIX::J_SADIN | Freedom isn't free. | Mon Jun 02 1997 20:04 | 7 |
|
found chiller heater was turned off. Please leave the heaters on.
Thanks.
a-shift
|
23.96 | primary table motor leaking oil | FABSIX::J_SADIN | Freedom isn't free. | Mon Jun 02 1997 20:06 | 9 |
|
Weekly pm completed. Found green oil from motor (primary table) in
catch pan. Bruce Tower is working the issue with Speedfam.
a-shift
|
23.97 | oil leakage on .b4 | FABSIX::J_SADIN | Freedom isn't free. | Wed Jun 04 1997 17:45 | 7 |
|
Checked oil leakage at Bruce Tower's request. Found about 1/2 cup
of oil in catch pan. Cleaned up what was there.
a-shift
|